OpenStudy (anonymous):

Opening day attendance at a particular ball field dropped from 45,000 people in 2001 to 31,000 people in 2006. Find the rate of change in opening day attendance.

OpenStudy (anonymous):

Okay so, 2006-2001=5 years. 45,000-31,000=14000 Rate= 14,000/5 years Rate=2,800 students/year.